Starting from rest, you push your 1000kg car over a 5m distance, on an horizontal ground, applying a horizontal 400N force.a) What is the car kinetic energy change?b) What is its final velocity at the end of the 5 meters displacement? Disregard any friction force.


1
Expert's answer
2021-01-31T03:41:32-0500

a) From the work-kinetic energy theorem, we get:


"\Delta KE=W=Fd=400\ N\cdot5\ m=2000\ J."

b)

"KE_f=\dfrac{1}{2}mv_f^2=W,""v_f=\sqrt{\dfrac{2W}{m}}=\sqrt{\dfrac{2\cdot2000\ J}{1000\ kg}}=2\ \dfrac{m}{s}."